A rectangle with height and width equal to 3 and 19 respectively, is drawn on a checkered paper. Bazil paints a random horizontal 1 x 2 rectangle, and Peter paints a random vertical 2 x 1 rectangle (each rectangle consists of 2 sells). Find the probability that at least one of the cells is painted twice. Express the answer in percent, and round to the nearest integer.


1
Expert's answer
2020-11-26T13:34:57-0500

We assume that Bazil and Peter paint on the drawn rectangle. The picture below presents

rectangle "5\times6".



As we can see, it is possible to paint 25 different horizontal "1\times2" rectangles. In a similar way we conclude that on rectangle "3\times 19" it is possible to paint "3 \cdot 18 =54" horizontal "1\times2" rectangles and "2 \cdot 19=38" different vertical "2\times1" rectangles. We have "2052" possible combinations in total.

On "2\times2" square it is possible to paint 4 different intersections of vertical and horizontal rectangles. We can consider "2 \cdot 18=36" different "2\times2" squares on the rectangle. Thus, we obtain "36 \cdot 4=144" different combination, when two rectangles intersect each other. I.e., there will be cells painted twice.

Thus, the probability that one cell is painted twice is: "p=\frac{144}{2052}=0.07 = 7\%"